Consider a call option with one year time remaining to expiration and the table below

(15 pts)

Time

Stock Price

Call Exercise Price

Call premium

Now

32

30

4

Later

29

30

1

      Provide the result of the covered call writing strategy in the table below in the context of marking to market only. Write as to what you are doing in each block and show numbers.

Time

Stock Position

Call Option Position

Net or combined position

Now

Action:

Action:

Result:

6 months later

Action:

Action:

Result:

Results in $

Results in %
Answer #1

Assumption: mentioned call option is a European option and can only be exercised at expiration (end of one year)

Time

Stock Position

Call Option Position

Net or combined position

Now

Action:

Buy Stock for $32

Therefore, outflow of $32

Action:

Enter into a short call position earning premium of $4

Therefore, cash inflow of $4

Result:

-32 + 4 = -$28

6 months later

Action:

Current market value of stock = $29

Action:

since option is not expired yet, no action

Result:

Mark to market value = +$29

Results in $

-28 + 29 = +1

Results in %

(1 / 28) * 100 = 3.57%